# InfiniBench: Infinite Benchmarking for Visual Spatial Reasoning with Customizable Scene Complexity

### v1

@@ 4.1 Experimental Setup
Implementation. We use Gemini-2.5-Pro comanici2025gemini in the agentic framework for scene constraint generation, and our layout optimizer is built upon Infinigen raistrick2024infinigen, leveraging its rich and robust asset library. For camera trajectory generation and final 3D rendering, we enforce a pipeline with Trimesh trimesh for geometric processing, PyRender pyrender for fast occlusion checks, and Blender’s Cycles engine blender for high-fidelity video output. More details are in Appendix F.
Baselines. We compare InfiniBench with two categories of 3D scene generation methods. 1) Procedural Frameworks including Infinigen raistrick2024infinigen (using its original optimizer) and Luminous zhao2021luminous; and 2) LLM-based Layout Generators including LayoutGPT feng2023layoutgpt, Holodeck yang2024holodeck and I-Design ccelen2024design, which directly generate scene layouts from text. For fair comparisons, the same 3D asset library (provided by Infinigen) is used in all experiments.
@@
  • Text-image alignment, which measure the semantic alignment as the CLIP score  radford2021learning between the text prompt and the rendered top-down image of generated scene.
  • Layout realism, scored by a GPT-5 evaluator as suggested by Scenethesis ling2025scenethesis, incorporating key factors such as spatial coherence and functional plausibility.
  • Physical plausibility, measured by the number of physically impossible artifacts, especially the count of out-of-boundary objects (#OB) and pairs of collided objects (#CN), as suggested by SceneWeaver yang2025sceneweaver.
@@ 4.2 Main Results
Quantitative Results. Tables 1 and 2 show that as the scene complexity increases, current generation methods exhibit distinct tradeoffs among different evaluation metrics. LLM-based generation methods reach the required complexity constraints and maintain relatively high Fidelity at the cost of lower physical plausibility, as evidenced by an increase in #OB and #CN. For instance, with a high amount of objects in the scene, LayoutGPT’s collision count rises to 13.5, 5.6x higher than that with a low amount of objects. Conversely, procedural generation frameworks excel at maintaining physical plausibility, but have difficulty in adhering to the prompt’s quantitative requirements of scene complexity. As such required scene complexity increases, their Fidelity scores drop drastically (e.g., 0.42 and 0.64 for Luminous and Infinigen, respectively), indicating a failure of generating the required number of objects.
@@ 5 Using InfiniBench to Diagnose VLMs’ Failures in Spatial Reasoning
To further verify the usefulness of InfiniBench in diagnosing VLMs’ abilities and failure modes in spatial reasoning with high complexity, we apply the benchmarks generated by InfiniBench with different levels of compositional, relational and observational complexities to the following set of representative spatial reasoning tasks brown2025sims. We examine where and why VLM’s spatial reasoning succeeds or fails, with two close-sourced models (Gemini-2.5-Pro comanici2025gemini, GPT-5 openai_gpt5_system_card) and three open-sourced models (LLaVa-Video-7B zhang2024videoinstructiontuningsynthetic, InternVL3.5-8B wang2025internvl3_5, Cambrian-S-7B yang2025cambrian).
Measurement tasks that probe VLM’s ability of reasoning about metric properties and object scale. We generate tasks that require both dimensional comparison and fine-grained distance estimation based on complex referring expressions, such as Q: What’s the [Height] of the [Object_Name] with the texture of [Material_name] and [Relation to a nearby object]. Performance will be measured by mean relative accuracy between numerical answer and ground truth yang2025thinking.
Perspective-taking tasks that assess compositional understanding from a given viewpoint. We choose object counting tasks with various numbers of irrelevant objects, such as Q: How many [Object_Name] are in the video. Performance will be similarly measured by mean relative accuracy yang2025thinking.
@@ 5.1 VLM’s Spatial Reasoning with Varying Compositional Scene Complexity
We constructed benchmarks with the number of objects from 5 to 50, as shown in Figure 9. Results in Figure 10 showed that the reasoning performance of VLMs declines markedly with more objects, revealing a shared difficulty in managing more complicated visual clutter. Gemini-2.5-Pro and Cambrian-S-7B achieve higher accuracy compared to other models, as the scene complexity grows. We hypothesize that the performance degradation with high compositional scene complexity stems from repetitive counting, since the same object could appears multiple times in the input video, an effect reflected in the models’ tendency to overestimate counts relative to the ground truth.
@@ Appendix A In-Context Learning for Scene Constraint Generation
In our design of InfiniBench, the LLM takes the following inputs: the user’s scene description in natural language, the given descriptions about the list of available assets and constraint APIs (determined by the specific procedural generation framework being used). Besides, to ensure the generated scene constraints are machine-readable and physically plausible, in-context examples are also provided to the LLM to demonstrate the specific syntax of the underlying constraint language used in the procedural generation framework. In practice, we use the default home constraints which are manually crafted as the in-context examples. Some more specific examples, with respect to the constraint APIs and syntax being used in the Infinigen raistrick2024infinigen framework, are provided below.
@@ A.1 Example 1: Dining Rooms
3diningtables = furniture[Semantics.Table][tables.TableDiningFactory]
4diningchairs = furniture[Semantics.Chair][seating.ChairFactory]
5constraints["dining_chairs"] = rooms.all(
@@
18score_terms["dining_chairs"] = rooms.all(
@@
29constraints["dining_table_objects"] = rooms.all(
@@
49score_terms["dining_table_objects"] = rooms.mean(
@@
59diningrooms = rooms[Semantics.DiningRoom].excludes(cu.room_types)
60constraints["diningroom"] = diningrooms.all(
@@
75)
76score_terms["diningroom"] = diningrooms.mean(
@@ A.2 Example 2: Bathrooms
1# region BATHROOMS
2bathrooms = rooms[Semantics.Bathroom].excludes(cu.room_types)
@@
4toilet = wallfurn[bathroom.ToiletFactory]
5bathtub = wallfurn[bathroom.BathtubFactory]
6sink = wallfurn[bathroom.StandingSinkFactory]
@@
8hardware = obj[bathroom.HardwareFactory].related_to(bathrooms, cu.against_wall)
@@
10constraints["bathroom"] = bathrooms.all(
@@
23score_terms["toilet"] = rooms.all(
@@
32constraints["bathtub"] = bathrooms.all(
@@
37)
38score_terms["bathtub"] = bathrooms.all(
@@
50score_terms["bathroom"] = (
@@
52) + cl.accessibility_cost(mirror, furniture, cu.down_dir).maximize(weight=3)
53# endregion
@@ Appendix B CoT Reasoning for Refinement of Scene Constraints
A detailed end-to-end example of such CoT reasoning, with the constraint APIs and syntax provided by the Infinigen framework raistrick2024infinigen, is provided in Figure 13 below:
@@
Figure 14: Examples of generated living rooms with varying levels of compositional complexity, measured by the number of objects. ‘Low’ indicates seating for fewer than five people; ‘Medium’ indicates seating for five to ten people; and ‘High’ indicates seating for more than ten people. The user’s prompt to the LLM agent is: “Generate a living room with seating for ** people.”
Figure 15: Examples of generated bedrooms with varying levels of compositional complexity, measured by the number of objects. ‘Low’ indicates only one shelf; ‘Medium’ indicates 2-8 shelves; and ‘High’ indicates seating for more than 10 shelves. The user’s prompt to the LLM agent is: “Generate a bedroom with ** shelves.”
Figure 16: Examples of generated dining rooms with varying levels of compositional complexity, measured by the number of objects. ‘Low’ indicates seating for fewer than five people; ‘Medium’ indicates seating for five to ten people; and ‘High’ indicates seating for more than ten people. The user’s prompt to the LLM agent is: “Generate a dining room with seating for ** people.”
@@ Appendix D Gallery of Generated 3D Scenes
We provide additional examples of the 3D scenes generated by InfiniBench, demonstrating its versatility across various room types and levels of scene complexity. The scenes with different levels of compositional complexity are shown in Figures 14-16, and the scenes with different levels of relational complexity are shown in Figures 17-19. In addition, in Figure 20 and Figure 21, we also provided examples of the optimized camera trajectories, along with the scene frames at the selected waypoints, with different levels of observational complexities. Note that, in addition to InfiniBench’s native scene-generation pipeline, we also leverage the room-shape generation module from Infinigen to produce more realistic renderings of these example scenes.
Figure 17: Examples of generated living rooms with varying levels of relational complexity, measured by the room occupancy ratio. ‘Low’ indicates an occupancy rate below 10%; ‘Medium’ indicates 10–50%; and ‘High’ indicates an occupancy rate above 50%. The user’s prompt to the LLM agent is: “Generate a living room with with occupancy rate of **.”
Figure 18: Examples of generated bedrooms with varying levels of relational complexity, measured by the room occupancy ratio. ‘Low’ indicates an occupancy rate below 10%; ‘Medium’ indicates 10–50%; and ‘High’ indicates an occupancy rate above 50%. The user’s prompt to the LLM agent is: “Generate a bedroom with with occupancy rate of **.”
Figure 19: Examples of generated dining rooms with varying levels of relational complexity, measured by the room occupancy ratio. ‘Low’ indicates an occupancy rate below 10%; ‘Medium’ indicates 10–50%; and ‘High’ indicates an occupancy rate above 50%. The user’s prompt to the LLM agent is: “Generate a dining room with with occupancy rate of **.”
Figure 20: Visualization of the camera trajectory under a specific level of observational complexity (partial occlusion). The user’s prompt to the LLM agent is: “Generate a typical living room. The camera should capture objects with partial occlusion.”
Figure 21: Visualization of the camera path under a specific level of observational complexity (bird-eye view, camera height is 2m). The user’s prompt to the LLM agent is: “Generate a typical living room. The camera should capture height should be 2m”

### v2

@@ 4.1 Experimental Setup
Implementation. We use Gemini-2.5-Pro [comanici2025gemini] in the agentic framework for scene constraint generation, and our layout optimizer is built upon Infinigen [raistrick2024infinigen], leveraging its rich and robust asset library. For camera trajectory generation and final 3D rendering, we enforce a pipeline with Trimesh [trimesh] for geometric processing, PyRender [pyrender] for fast occlusion checks, and Blender’s Cycles engine [blender] for high-fidelity video output. More details are in Appendix F.
Baselines. We compare InfiniBench with two categories of 3D scene generation methods. 1) Procedural Frameworks including Infinigen [raistrick2024infinigen] (using its original optimizer) and Luminous [zhao2021luminous]; and 2) LLM-based Layout Generators including LayoutGPT [feng2023layoutgpt], Holodeck [yang2024holodeck] and I-Design [ccelen2024design], which directly generate scene layouts from text. For fair comparisons, the same 3D asset library (provided by Infinigen) is used in all experiments.
@@
  • Text-image alignment, which measure the semantic alignment as the CLIP score  [radford2021learning] between the text prompt and the rendered top-down image of generated scene.
  • Layout realism, scored by a GPT-5 evaluator as suggested by Scenethesis [ling2025scenethesis], incorporating key factors such as spatial coherence and functional plausibility.
  • Physical plausibility, measured by the number of physically impossible artifacts, especially the count of out-of-boundary objects (#OB) and pairs of collided objects (#CN), as suggested by SceneWeaver [yang2025sceneweaver].
@@ 4.2 Main Results
Quantitative Results. Tables 1 and 2 show that as the scene complexity increases, current generation methods explicit distinct tradeoffs among different evaluation metrics. LLM-based generation methods reach the required complexity constraints and maintain relatively high Fidelity at the cost of lower physical plausibility, as evidenced by an increase in #OB and #CN. For instance, with a high amount of objects in the scene, LayoutGPT’s collision count rises to 13.5, 5.6x higher than that with a low amount of objects. Conversely, procedural generation frameworks excel at maintaining physical plausibility, but have difficulty in adhering to the prompt’s quantitative requirements of scene complexity. As such required scene complexity increases, their Fidelity scores drop drastically (e.g., 0.42 and 0.64 for Luminous and Infinigen, respectively), indicating a failure of generating the required number of objects.
@@ 5 Using InfiniBench to Diagnose VLMs’ Failures in Spatial Reasoning
To further verify the usefulness of InfiniBench in diagnosing VLMs’ abilities and failure modes in spatial reasoning with high complexity, we apply the benchmarks generated by InfiniBench with different levels of compositional, relational and observational complexities to the following set of representative spatial reasoning tasks [brown2025sims]. We examine where and why VLM’s spatial reasoning succeeds or fails, with two close-sourced models (Gemini-2.5-Pro [comanici2025gemini], GPT-5 [openai_gpt5_system_card]) and three open-sourced models (LLaVa-Video-7B [zhang2024videoinstructiontuningsynthetic], InternVL3.5-8B [wang2025internvl3_5], Cambrian-S-7B [yang2025cambrian]).
Measurement tasks that probe VLM’s ability of reasoning about metric properties and object scale. We generate tasks that require both dimensional comparison and fine-grained distance estimation based on complex referring expressions, such as Q: What’s the [Height] of the [Object_Name] with the texture of [Material_name] and [Relation to a nearby object]. Performance will be measured by mean relative accuracy between numerical answer and ground truth [yang2025thinking].
Perspective-taking tasks that assess compositional understanding from a given viewpoint. We choose object counting tasks with various numbers of irrelevant objects, such as Q: How many [Object_Name] are in the video. Performance will be similarly measured by mean relative accuracy [yang2025thinking].
@@ 5.1 VLM’s Spatial Reasoning with Varying Compositional Scene Complexity
We constructed benchmarks with the number of objects from 5 to 50, as shown in Figure 9. Results in Figure 10 showed that the reasoning performance of VLMs declines markedly with more objects, revealing a shared difficulty in managing more complicated visual clutter. Gemini-2.5-Pro and Cambrian-S-7B achieve higher accuracy compared to other models, as the scene complexity grows. We hypotheisze that the performance degradation with high compositional scene complexity stems from repetitive counting, since the same object could appears multiple times in the input video, an effect reflected in the models’ tendency to overestimate counts relative to the ground truth.
@@ Appendix A In-Context Learning for Scene Constraint Generation
In our design of InfiniBench, the LLM takes the following inputs: the user’s scene description in natural language, the given descriptions about the list of available assets and constraint APIs (determined by the specific procedural generation framework being used). Besides, to ensure the generated scene constraints are machine-readable and physically plausible, in-context examples are also provided to the LLM to demonstrate the specific syntax of the underlying constraint language used in the procedural generation framework. In practice, we use the default home constraints which are manually crafted as the in-context examples. Some more specific examples, with respect to the constraint APIs and syntax being used in the Infinigen [raistrick2024infinigen] framework, are provided below.
@@ A.1 Example 1: Dining Rooms
3 diningtables = furniture[Semantics.Table][tables.TableDiningFactory]
4 diningchairs = furniture[Semantics.Chair][seating.ChairFactory]
5 constraints["dining_chairs"] = rooms.all(
@@
18 score_terms["dining_chairs"] = rooms.all(
@@
29 constraints["dining_table_objects"] = rooms.all(
@@
49 score_terms["dining_table_objects"] = rooms.mean(
@@
59 diningrooms = rooms[Semantics.DiningRoom].excludes(cu.room_types)
60 constraints["diningroom"] = diningrooms.all(
@@
75 )
76 score_terms["diningroom"] = diningrooms.mean(
@@ A.2 Example 2: Bathrooms
1 # region BATHROOMS
2 bathrooms = rooms[Semantics.Bathroom].excludes(cu.room_types)
@@
4 toilet = wallfurn[bathroom.ToiletFactory]
5 bathtub = wallfurn[bathroom.BathtubFactory]
6 sink = wallfurn[bathroom.StandingSinkFactory]
@@
8 hardware = obj[bathroom.HardwareFactory].related_to(bathrooms, cu.against_wall)
@@
10 constraints["bathroom"] = bathrooms.all(
@@
23 score_terms["toilet"] = rooms.all(
@@
32 constraints["bathtub"] = bathrooms.all(
@@
37 )
38 score_terms["bathtub"] = bathrooms.all(
@@
50 score_terms["bathroom"] = (
@@
52 ) + cl.accessibility_cost(mirror, furniture, cu.down_dir).maximize(weight=3)
53 # endregion
@@ Appendix B CoT Reasoning for Refinement of Scene Constraints
A detailed end-to-end example of such CoT reasoning, with the constraint APIs and syntax provided by the Infinigen framework [raistrick2024infinigen], is provided in Figure 13 below:
@@
Figure 14: Examples of generated living rooms with varying levels of compositional complexity, measured by the number of objects. ‘Low’ indicates seating for fewer than five people; ‘Medium’ indicates seating for five to ten people; and ‘High’ indicates seating for more than ten people. The input prompt to the LLM agent for scene generation is: “Generate a living room with seating for <<num>> people”.
Figure 15: Examples of generated bedrooms with varying levels of compositional complexity, measured by the number of objects. ‘Low’ indicates only one shelf; ‘Medium’ indicates 2-8 shelves; and ‘High’ indicates seating for more than 10 shelves. The input prompt to the LLM agent for scene generation is: “Generate a bedroom with <<num>> shelves”.
Figure 16: Examples of generated dining rooms with varying levels of compositional complexity, measured by the number of objects. ‘Low’ indicates seating for fewer than five people; ‘Medium’ indicates seating for five to ten people; and ‘High’ indicates seating for more than ten people. The input prompt to the LLM agent for scene generation is: “Generate a dining room with seating for <<num>> people”.
@@ Appendix D Gallery of Generated 3D Scenes
We provide additional examples of the 3D scenes generated by InfiniBench, demonstrating its versatility across various room types and levels of scene complexity. The scenes with different levels of compositional complexity are shown in Figures 14-16, and the scenes with different levels of relational complexity are shown in Figures 17-19. For each generated scene, we showed images captured from two viewpoints, for better visualization of the generated room layouts and furniture positioning. In addition, in Figure 20 and Figure 21, we also provided examples of the optimized camera trajectories, along with the scene frames at the selected waypoints, with different levels of observational complexities. Note that, in addition to InfiniBench’s native scene-generation pipeline, we also leverage the room-shape generation module from Infinigen to produce more realistic renderings of these example scenes.
Figure 17: Examples of generated living rooms with varying levels of relational complexity, measured by the room occupancy ratio. ‘Low’ indicates an occupancy ratio below 10%; ‘Medium’ indicates an occupancy ratio between 10% and 50%; and ‘High’ indicates an occupancy rate above 50%. The input prompt to the LLM agent for scene generation is: “Generate a living room with a room occupancy ratio of <<ratio>>”.
Figure 18: Examples of generated bedrooms with varying levels of relational complexity, measured by the room occupancy ratio. ‘Low’ indicates an occupancy ratio below 10%; ‘Medium’ indicates an occupancy ratio between 10% and 50%; and ‘High’ indicates an occupancy rate above 50%. The input prompt to the LLM agent for scene generation is: “Generate a bedroom with a room occupancy ratio of <<ratio>>”.
Figure 19: Examples of generated dining rooms with varying levels of relational complexity, measured by the room occupancy ratio. ‘Low’ indicates an occupancy ratio below 10%; ‘Medium’ indicates an occupancy ratio between 10% and 50%; and ‘High’ indicates an occupancy rate above 50%. The input prompt to the LLM agent for scene generation is: “Generate a dining room with a room occupancy ratio of <<ratio>>”.
Figure 20: Visualization of the camera trajectory under a specific level of observational complexity (partial occlusion). The input prompt to the LLM agent for scene generation is: “Generate a typical living room. The camera should capture objects with partial occlusion”.
Figure 21: Visualization of the camera path under a specific level of observational complexity (bird-eye view, camera height is 2m). The input prompt to the LLM agent for scene generation is: “Generate a typical living room. The camera’s capture height should be 2 meters”.